one half of the total of 6 times a number and 20 is 4. what is the number?
thanks...

Answer by checkley77(12569)

(6X+20)/2=4 CROSS MULTIPLY
6X+20=8
6X=8-20
6X=-12
X=-12/6
X=-2 ANSWER.
PROOF:
(6*-2+20)/2=4
(12-20)/2=4
8/2=4
4=4